What is the pH change when 19.5 mL of 0.108 M NaOH Is added to 71.0 mL of a buffer solution consisting of 0.127 M NH3 and 0.161 M NH4? ( for ammonium ion is 5.6*10^-10.) pH change =
Answer #1

ka for NHy + = 5.68 10-10 Kb for NH3 = 10-14 kw 10-14 10-14 = 1.79 x 10-5 ka 5.68 10-10 Befor adding NaOH [NH3 ] = 0.127M [NHICE Table + + H₂O(1) concentration (MI NHy + (aq) 0.12631 - 0.02327 OH lag) 0.02327 - 0.02327 NH3199) 0.09964 + 0.02327 0.122